OpenStudy (anonymous):

If the equation of a circle is (x + 4)2 + (y - 6)2 = 25, its center point is (4, 6) (-4, 6) (4, -6)

OpenStudy (alexandervonhumboldt2):

(x-a)^2+(y-b)^2=r^2 where (a, b) is the center

OpenStudy (solomonzelman):

\(\LARGE\color{slate}{ (x -\color{red}{h})^2 + (y - \color{green}{k})^2 = \color{blue}{r}^2 }\) is the formula. where: \(\LARGE\color{slate}{ (\color{red}{h},\color{green}{k}) }\) is the center. \(\LARGE\color{slate}{\color{blue}{r} }\) is the radius.

OpenStudy (anonymous):

so -4,6?

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (solomonzelman):

yes, (-4,6) is correct.

OpenStudy (anonymous):

thanks:))

OpenStudy (solomonzelman):

yw